OpenStudy (anonymous):

When you multiply 2 polynomials that have an exponent, do you add the exponents or multiply them?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

Add is correct. X^2 * x^3 = x^5 BUT (x^2)^3 = x^6

OpenStudy (anonymous):

Thank you!